# Graph Flow Matching: Enhancing Image Generation with Neighbor-Aware Flow Fields

### v2

@@ Abstract
Flow matching casts sample generation as learning a continuous-time velocity field that transports noise to data. Existing flow matching networks typically predict each point’s velocity independently, considering only its location and time along its flow trajectory, and ignoring neighboring points. However, this pointwise approach may overlook correlations between points along the generation trajectory that could enhance velocity predictions, thereby improving downstream generation quality. To address this, we propose Graph Flow Matching (GFM), a lightweight enhancement that decomposes the learned velocity into a reaction term—any standard flow matching network—and a diffusion term that aggregates neighbor information via a graph neural module. This reaction–diffusion formulation retains the scalability of deep flow models while enriching velocity predictions with local context, all at minimal additional computational cost. Operating in the latent space of a pretrained variational autoencoder, GFM consistently improves Fréchet Inception Distance (FID) and recall across five image generation benchmarks (LSUN Church, LSUN Bedroom, FFHQ, AFHQ-Cat, and CelebA-HQ at $256\times 256$), demonstrating its effectiveness as a modular enhancement to existing flow matching architectures.
@@ 1 Introduction
Flow matching has recently gained traction as a promising generative modeling paradigm, framing sample synthesis as the integration of a learned (interpolated) continuous-time velocity field that deterministically or stochastically transforms noise into data lipman2023flowmatchinggenerativemodeling ; chen2018neural . Given two distributions $\pi_{0}$ and $\pi_{1}$, the task is to learn a velocity field $v(\mathbf{x},t)$ that, when integrated, solves:
$\dot{\bf x}=v(\mathbf{x},t),\quad\mathbf{x}(0)\sim\pi_{0},\quad\mathbf{x}(1)\sim\pi_{1}.$ (1)
While highly effective, current flow matching architectures predict each point’s velocity independently, conditioning only on its $(\mathbf{x},t)$ coordinates without considering neighboring points. This overlooks a fundamental property of high-dimensional data: local coherence, where nearby points typically exhibit similar behaviors and follow in similar trajectories.
@@
We propose Graph Flow Matching (GFM), which enhances standard flow networks with neighbor awareness through a reaction-diffusion decomposition:
$v(\mathbf{x},t)=v_{\text{react}}(\mathbf{x},t)+v_{\text{diff}}(\mathbf{x},t;\,\mathcal{N}(\mathbf{x},t)),$ (2)
@@
where $v_{\text{react}}$ is any standard flow network, and $v_{\text{diff}}$ is a lightweight graph-based correction term that aggregates information from neighboring samples $\mathcal{N}(\mathbf{x},t)$. Figure 1 illustrates this process, where the flow trajectory from $\pi_{0}$ to $\pi_{1}$ is augmented by incorporating neighborhood structure via graphs at intermediate time steps. Designing a flow matching pipeline requires two key decisions: network architecture and training strategy (e.g., Rectified Flow RectifiedFlow_liu2022flow , Consistency FM ConsistencyFM_yang2024consistencyfm ). We focus on the architecture aspect, proposing an enhancement compatible with any training objective and existing base flow network - pretrained or otherwise.
@@
In this work, we adopt the latent-space generative modeling paradigm, popularized by work such as Stable Diffusion StableDiffusion_rombach2022high , and recently used for flow matching as in LFM dao2023flow , where image generation is performed in a perceptually compressed latent space produced by a pretrained variational autoencoder (VAE) VAE_kingma2022autoencodingvariationalbayes . This choice reduces computational cost as the latent space defines meaningful distance metrics for constructing graphs, which are easy to compute. Our contributions therefore operate entirely within this latent framework. Figures 23 qualitatively demonstrate our method’s effectiveness across three diverse datasets. Quantitatively, we validate GFM on five unconditional image generation benchmarks—LSUN Church LSUN_dataset_yu15lsun , LSUN Bedroom LSUN_dataset_yu15lsun , FFHQ FFHQ_dataset_karras2019style , AFHQ-Cat AFHQ_dataset_choi2020starganv2 , and CelebA-HQ CelebAHQ_dataset_karras2018progressive at $256\times 256$ resolution, showing consistent improvements in Fréchet Inception Distance (FID) and recall while adding minimal computational overhead ($\lesssim 10\%$ additional parameters).
@@
(i) We introduce Graph Flow Matching (GFM), a modular reaction–diffusion framework that enhances flow networks with neighbor-aware velocity correction. GFM decomposes the velocity field into a standard pointwise term (using any existing flow network) and a graph-based diffusion term, integrating seamlessly with existing backbones and training strategies without requiring modifications to losses or solvers; (ii) We validate GFM in the latent space of a pretrained VAE, demonstrating consistent improvements in unconditional image generation across five standard benchmarks, establishing our approach as a practical enhancement to existing flow matching architectures.
Figure 1: Graph flow matching enriches the flow trajectory from the initial distribution ($t=0$) to the target distribution ($t=1$) by incorporating neighborhood information at intermediate time steps. Graphs at each step—constructed via attention—guide local interactions; dashed edges indicate weaker influences.
Figure 2: Neighbor-aware flow matching enhances image generation. FFHQ samples ($256\times 256$) generated using the same random seed by: (top) baseline ADM U-Net dao2023flow , (middle) ADM with MPNN-based correction, and (bottom) ADM with a GPS-based correction module GPS_rampavsek2022recipe . GFM variants produce more coherent facial features and sharper details compared to the baseline model.
@@ 2 Background and Related Work
A central challenge in generative modeling is transporting a simple base distribution (e.g., Gaussian noise) into a complex, high-dimensional target distribution. This underpins models such as generative adversarial networks (GANs) GAN_goodfellow2014generative , variational autoencoders (VAEs) VAE_kingma2022autoencodingvariationalbayes , normalizing flows papamakarios_NormalizingFlows , and diffusion models song2019generative ; song2021scorebasedgenerativemodelingstochastic ; DDPM_ho2020denoising . Recent trends increasingly adopt continuous-time formulations. Continuous normalizing flows (CNFs) chen2018neural generalize discrete flows by learning transformations via ODEs. Diffusion models, originally discrete-time DDPM_ho2020denoising , have evolved into continuous-time score-based models defined through stochastic differential equations (SDEs) song2021scorebasedgenerativemodelingstochastic . Flow matching lipman2023flowmatchinggenerativemodeling offers a deterministic, continuous-time alternative that avoids score estimation by directly supervising a velocity field that transports noise to data. It combines fast sampling and simplified training, making it an attractive framework for generative modeling. Our work builds upon this foundation by introducing a graph-based architectural enhancement that enriches velocity predictions with information from neighboring points along the flow trajectory. On a different note, while recent works use flow matching for graph generation VarFM_eijkelboom2024variational ; song2023equivariant_3DMoleculeGeneration ; Defog_qin2024defog ; GraphCondFlowMatching_scassola2025graph ; LatentSpaceGraphDiffusion_pombala2025exploring , our GFM approach focuses on a different goal; it uses graph learning techniques for enhancing flow matching techniques.
@@ 2.1 Flow Matching Preliminaries
Flow matching formulates sample generation as solving an ordinary differential equation (ODE) that transports a source distribution $\pi_{0}$ to a target distribution $\pi_{1}$ over a finite time horizon $t\in[0,1]$. Given two random variables $X_{0}\sim\pi_{0}$ and $X_{1}\sim\pi_{1}$ with support in $\mathcal{X}\subseteq\mathbb{R}^{d}$, the goal is to learn a velocity field $v_{\theta}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ such that the solution $\mathbf{x}(t)$ to the initial value problem
$\dot{\bf x}=v_{\theta}(\mathbf{x},t),\quad\mathbf{x}(0)\sim\pi_{0},$ (3)
@@
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{1}}\left[\left\|v_{\theta}\left(\mathbf{x}_{t},t\right)-\mathbf{v}^{*}(\mathbf{x}_{t},t)\right\|^{2}\right].$ (5)
@@
2.2 Limitations of Pointwise Velocity Prediction
Current flow matching architectures predominantly adopt a pointwise prediction paradigm, directly regressing from query coordinates $(\mathbf{x}_{t},t)$ to velocity $v(\mathbf{x}_{t},t)$ while treating each point as an independent interpolation task. This mirrors the limitation of early point cloud architectures like PointNet (qi2017pointnet, ), which process points independently to interpolate features and thus may struggle to model local geometric relationships, a constraint that PointNet++ (qi2017pointnetplusplus, ) directly addressed by introducing hierarchical neighborhood aggregation. Similarly, methods like DGCNN (DGCNN_paper, ) introduced dynamic graph construction, showing that explicitly modeling local neighborhood structure improves performance over purely pointwise processing.
Without such neighborhood awareness, pointwise methods may struggle in regions with sharp transitions, sparse supervision, or multiple modes popeintrinsic_dim_of_images ; wanggeometric_analysis , and pointwise interpolation may fail to recover the underlying transport dynamics, leading to artifacts or degraded sample quality.
The core issue stems from overlooking a crucial inductive bias: local coherence (localCoherence1_wang_6064953, ). In high-dimensional manifolds, points that are spatially proximate often belong to the same semantic region and follow similar trajectories (ManifoldLearning_MUNOZ2023115859, ; PointOT_9763535, ). Just as nearby points in classification tasks tend to share labels, neighboring interpolation points in flow matching—particularly those along a shared trajectory—typically exhibit similar velocities. Neglecting this correlation structure may limit generalization, especially in regions of high curvature or sparse sampling where models must interpolate with limited local support (Lee_math11102375, ).
@@ 2.3 Graph-based Approaches in Deep Learning
The limitations of pointwise processing have motivated significant research into graph-based architectures that explicitly model relationships between data points. Graph Neural Networks (GNNs) have emerged as a powerful framework for incorporating local structure, enabling each point to aggregate information from its neighbors through message passing mechanisms GCN_paper_kipf2017semi ; GAT_paper_velickovic2018graph .
In geometric deep learning, Graph Convolutional Networks (GCNs) GCN_paper_kipf2017semi , Graph Attention Networks (GATs) GAT_paper_velickovic2018graph , and Message Passing Neural Networks (MPNNs) MPNN_paper_gilmer2017neural have demonstrated the value of neighborhood aggregation across diverse domains. These architectures naturally encode inductive biases about local structure, leading to improved generalization in tasks ranging from molecular property prediction to social network analysis.
Recent work has also explored the connection between GNNs and partial differential equations (PDEs). PDE_GCN_NEURIPS2021_1f9f9d8f showed that the dynamics of PDEs can be encoded into GNN architectures, with the implicit PDE structure greatly influencing their suitability for specific tasks. This PDE-GNN connection provides a principled framework for designing graph architectures that capture desired physical dynamics, which is a key inspiration for the MPNN graph correction module used in our experiments, as outlined in Section 4.
@@ 2.4 Physical Inspiration: Reaction-Diffusion Systems
Reaction-diffusion systems (turing1990chemical, ), which model how substances spread and interact in space, provide a natural framework for incorporating local interactions into flow matching. In these systems, the evolution of a quantity $u(\mathbf{x},t)$ is governed by:
@@
This decomposition inspires our approach: by treating the standard flow matching velocity network as a reaction term and adding a graph based diffusion term, we enable velocity predictions to adapt based on local flow patterns. In particular, the reaction-diffusion equation is known to generate complex patterns turing1990chemical that can explain many natural processes. Previous work has successfully utilized this equation in the context of GNNs haber2019imexnet ; choi2023gread ; eliasof2024graph for solving graph related problems.

### v3

@@ Abstract
Flow matching casts sample generation as learning a continuous-time velocity field that transports noise to data. Existing flow matching networks typically predict each point’s velocity independently, considering only its location and time along its flow trajectory, and ignoring neighboring points. However, this pointwise approach may overlook correlations with points along the generation trajectory that could enhance velocity predictions, thereby improving downstream generation quality. To address this, we propose Graph Flow Matching (GFM), a lightweight enhancement that decomposes the learned velocity into a reaction term – any standard flow matching network – and a diffusion term that aggregates neighbor information via a graph neural module. This reaction-diffusion formulation retains the scalability of deep flow models while enriching velocity predictions with local context, all at minimal additional computational cost. Operating in the latent space of a pretrained variational autoencoder, GFM consistently improves Fréchet Inception Distance (FID) and recall across five image generation benchmarks (LSUN Church, LSUN Bedroom, FFHQ, AFHQ-Cat, and CelebA-HQ at $256\times 256$), demonstrating its effectiveness as a modular enhancement to existing flow matching architectures.
@@ 1 Introduction
Flow matching has recently gained traction as a promising generative modeling paradigm, framing sample synthesis as the integration of a learned (interpolated) continuous-time velocity field that deterministically or stochastically transforms noise into data (lipman2023flowmatchinggenerativemodeling; chen2018neural). Given two distributions $\pi_{0}$ and $\pi_{1}$, the task is to learn a velocity field $\mathbf{v}(\mathbf{x},t)$ that, when integrated, solves:
$\dot{\bf x}=\mathbf{v}(\mathbf{x},t),\quad\mathbf{x}(0)\sim\pi_{0},\quad\mathbf{x}(1)\sim\pi_{1}.$ (1)
We may view learning the field $\mathbf{v}(\mathbf{x},t)$ as an interpolation problem over the joint domain of $\mathbf{x}$ (namely $\mathcal{X}\subseteq\mathbb{R}^{d}$), and $t$ (namely $[0,1]$): $\mathcal{X}\times[0,1]$, where we supervise the model at discrete $(\mathbf{x},t)$ pairs sampled from a chosen transport plan, and require it to generalize to a velocity field everywhere in $\mathcal{X}$. Given $\mathbf{x}(0)$, which is typically a sample from a Gaussian distribution, the path in $\mathcal{X}\times[0,1]$ obtained by solving Equation 1 is what we will term a trajectory. In the context of image generation, each $\mathbf{point}$ $\mathbf{x}(t)$ along such a trajectory is an image. While highly effective, current flow matching architectures predict each point’s velocity pointwise, conditioning only on its $(\mathbf{x},t)$ coordinates without considering neighboring points in adjacent trajectories. This overlooks an important structural property shown in the literature (gao2024gaussian; gao2024convergence; fukumizu2024flow) but not explicitly leveraged: the smoothness of $\mathbf{v}(\mathbf{x},t)$ over $\mathcal{X}\times[0,1]$. We observe that this regularity implies that neighboring points along nearby trajectories typically exhibit correlated behavior, which could be harnessed to improve flow estimation.
@@
In this paper, we leverage this observation, and propose Graph Flow Matching (GFM), which enhances standard flow networks with neighbor awareness through a reaction-diffusion decomposition:
$\mathbf{v}(\mathbf{x},t)=\mathbf{v}_{\text{react}}(\mathbf{x},t)+\mathbf{v}_{\text{diff}}(\mathbf{x},t;\,\mathcal{N}(\mathbf{x},t)),$ (2)
@@
where $\mathbf{v}_{\text{react}}$ is any standard flow network, and $\mathbf{v}_{\text{diff}}$ is a lightweight graph-based correction term (a graph neural network) that aggregates information from the set of neighbors of $\mathbf{x}$ at time $t$, denoted by $\mathcal{N}(\mathbf{x},t)$. Figure 1 illustrates this process, where the flow trajectory from $\pi_{0}$ to $\pi_{1}$ is augmented by incorporating neighborhood structure via graphs at intermediate time steps. Designing a flow matching pipeline requires two key decisions: network architecture and training strategy (e.g., Rectified Flow (RectifiedFlow_liu2022flow), Consistency FM (ConsistencyFM_yang2024consistencyfm)). We focus on the architecture aspect, proposing an enhancement compatible with any training strategy and existing base flow network $\mathbf{v}_{\text{react}}$, pretrained or otherwise.
@@
In this work, we adopt the latent-space generative modeling paradigm, popularized by work such as Stable Diffusion (StableDiffusion_rombach2022high), and recently used for flow matching as in LFM (dao2023flow), where image generation is performed in a perceptually compressed latent space produced by a pretrained variational autoencoder (VAE) (VAE_kingma2022autoencodingvariationalbayes). This choice reduces computational cost as the latent space defines meaningful distance metrics for constructing graphs, which are easy to compute. Our contributions therefore operate entirely within this latent framework. Therefore, in this work, each “point” $\mathbf{x}$ is the VAE latent encoding $\mathbf{x}\in\mathbb{R}^{4\times 32\times 32}$ of an image (in this paper, the latent encoding of a $256\times 256$ RGB image); all graphs operate over these latent encodings (otherwise known as latent codes). We shall denote these latent codes as $\mathbf{x}$ in this paper. Figure 2 and Figure 3 (in the Appendix) qualitatively demonstrate our method’s effectiveness across three diverse datasets. Quantitatively, we validate GFM on five unconditional image generation benchmarks—LSUN Church (LSUN_dataset_yu15lsun), LSUN Bedroom (LSUN_dataset_yu15lsun), FFHQ (FFHQ_dataset_karras2019style), AFHQ-Cat (AFHQ_dataset_choi2020starganv2), and CelebA-HQ (CelebAHQ_dataset_karras2018progressive) at $256\times 256$ resolution, showing consistent improvements in Fréchet Inception Distance (FID) and recall while adding minimal computational overhead ($\lesssim 10\%$ additional parameters).
@@
(i) We introduce Graph Flow Matching (GFM), a modular reaction–diffusion framework that enhances flow networks with neighbor-aware velocity correction. GFM decomposes the velocity field into a standard pointwise term (using any existing flow network) and a graph-based diffusion term, integrating seamlessly with existing backbones and training strategies without requiring modifications to losses or solvers; (ii) We validate GFM in the latent space of a pretrained VAE, demonstrating consistent improvements in unconditional image generation across five standard benchmarks. We use two architectures for $\mathbf{v}_{\text{diff}}$, which we call MPNN and GPS, to empirically validate that the use of a graph module in the flow model this way is a promising enhancement to existing flow architectures.
Figure 1: Graph flow matching enriches the flow trajectory from the initial distribution ($t=0$) to the target distribution ($t=1$) by connecting, at each intermediate time $t$ (shown as slices) nodes $\mathbf{x}$ (shown as dots) that are latent vectors (VAE codes) of distinct images using attention‑based similarity. The flow network output for each node $\mathbf{x}$ is its flow velocity $\mathbf{v}(\mathbf{x},t)$. Each dot (node) at $t=0$ represents a Gaussian noise image, while each dot (node) at $t=1$ represents a generated image. Dashed edges indicate lower attention weights.
Figure 2: Neighbor-aware flow matching enhances image generation. FFHQ samples ($256\times 256$) generated using the same random seed by: (top) baseline ADM U-Net (dao2023flow), (middle) ADM with MPNN-based correction, and (bottom) ADM with a GPS-based correction module (GPS_rampavsek2022recipe). GFM variants produce more coherent facial features and sharper details compared to the baseline model.
@@ 2 Background and Related Work
A central challenge in generative modeling is transporting a simple base distribution (e.g., Gaussian noise) into a complex, high-dimensional target distribution. This underpins models such as generative adversarial networks (GANs) (GAN_goodfellow2014generative), variational autoencoders (VAEs) (VAE_kingma2022autoencodingvariationalbayes), normalizing flows (papamakarios_NormalizingFlows), and diffusion models (song2019generative; song2021scorebasedgenerativemodelingstochastic; DDPM_ho2020denoising). Recent trends increasingly adopt continuous-time formulations. Continuous normalizing flows (CNFs) (chen2018neural) for example, generalize discrete-time flows by learning transformations via ODEs. Diffusion models, originally discrete-time (DDPM_ho2020denoising), have evolved into continuous-time score-based models defined through stochastic differential equations (SDEs) (song2021scorebasedgenerativemodelingstochastic). Flow matching (lipman2023flowmatchinggenerativemodeling) offers a deterministic, continuous-time alternative that avoids score estimation by directly supervising a velocity field that transports noise samples (that is, samples from a Gaussian distribution) to data. It combines fast sampling and simplified training, making it an attractive framework for generative modeling. Our work builds upon this foundation by introducing a graph-based architectural enhancement that enriches velocity predictions with information from neighboring points along the flow trajectory. On a different note, while recent works use flow matching for graph generation (VarFM_eijkelboom2024variational; song2023equivariant_3DMoleculeGeneration; Defog_qin2024defog; GraphCondFlowMatching_scassola2025graph; LatentSpaceGraphDiffusion_pombala2025exploring), our GFM approach focuses on a different goal; it uses graph learning techniques for enhancing flow matching techniques.
@@ 2.1 Flow Matching Preliminaries
Flow matching formulates sample generation as solving an ordinary differential equation (ODE) that transports a source distribution $\pi_{0}$ to a target distribution $\pi_{1}$ over a finite time horizon $t\in[0,1]$. Given two random variables $X_{0}\sim\pi_{0}$ and $X_{1}\sim\pi_{1}$ with support in $\mathcal{X}\subseteq\mathbb{R}^{d}$, the goal is to learn a velocity field $\mathbf{v}_{\theta}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ such that the solution $\mathbf{x}(t)$ to the initial value problem
$\dot{\bf x}=\mathbf{v}_{\theta}(\mathbf{x},t),\quad\mathbf{x}(0)\sim\pi_{0},$ (3)
@@
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{1}}\left[\left\|\mathbf{v}_{\theta}\left(\mathbf{x}_{t},t\right)-\mathbf{v}^{*}(\mathbf{x}_{t},t)\right\|^{2}\right].$ (5)
@@
2.2 Interpolation Perspective
The flow velocity field $\mathbf{v}(\mathbf{x}_{t},t)$ is smooth, typically Lipschitz continuous in $\mathbf{x}$ and $t$ (gao2024gaussian; gao2024convergence). Moreover, the learned field satisfies the continuity equation, ensuring conservative mass transport (albergo2022building; lipman2023flowmatchinggenerativemodeling). Learning $\mathbf{v}$ from a finite set of samples thus amounts to a scattered‐data interpolation problem: one observes velocities at sparse $(\mathbf{x}_{t},t)$ locations and seeks to reconstruct a smooth vector field throughout the domain. Classical interpolants, such as radial basis functions, splines, and moving least squares, recover smooth functions by fitting to nearby samples with local support (micchelli1984interpolation; de1978practicalSplines; levin1998_MLSapproximation).
By contrast, most flow networks predict each $\mathbf{v}(\mathbf{x}_{t},t)$ independently, ignoring correlations among neighboring points along adjacent flow trajectories. To explicitly instill this inductive bias into the architecture, we build an attention‐weighted graph over the batch at each timestep and use a lightweight GNN to allow each velocity prediction to draw on its neighbors, mimicking how interpolation stencils blend local information. We view the GNN term, $\mathbf{v}_{\text{diff}}$, as a learnable analogue to classical interpolation kernels, generalizing fixed support functions to trainable networks. Further discussion is included in Section A.1.
@@ 2.3 Graph-based Approaches in Deep Learning
The limitations of pointwise processing have motivated significant research into graph-based architectures that explicitly model relationships between data points. Graph Neural Networks (GNNs) have emerged as a powerful framework for incorporating local structure, enabling each point to aggregate information from its neighbors through message passing mechanisms (GCN_paper_kipf2017semi; GAT_paper_velickovic2018graph).
In geometric deep learning, Graph Convolutional Networks (GCNs) (GCN_paper_kipf2017semi), Graph Attention Networks (GATs) (GAT_paper_velickovic2018graph), and Message Passing Neural Networks (MPNNs) (MPNN_paper_gilmer2017neural) have demonstrated the value of neighborhood aggregation across diverse domains. These architectures naturally encode inductive biases about local structure, leading to improved generalization in tasks ranging from molecular property prediction to social network analysis.
Recent work has also explored the connection between GNNs and partial differential equations (PDEs). (PDE_GCN_NEURIPS2021_1f9f9d8f) showed that the dynamics of PDEs can be encoded into GNN architectures, with the implicit PDE structure greatly influencing their suitability for specific tasks. This PDE-GNN connection provides a principled framework for designing graph architectures that capture desired physical dynamics, which is a key inspiration for the MPNN graph correction module used in our experiments, as outlined in Section 4.
@@ 2.4 Physical Inspiration: Reaction-Diffusion Systems
Reaction-diffusion systems (turing1990chemical), which model how substances spread and interact in space, provide a natural framework for incorporating local interactions into flow matching. In these systems, the evolution of a quantity $u(\mathbf{x},t)$ is governed by:
@@
This decomposition inspires our approach: by treating the standard flow matching velocity network as a reaction term that captures pointwise dynamics and adding a graph based diffusion term that allows neighbor to neighbor communication, we enable velocity predictions to adapt based on local flow patterns. In particular, the reaction-diffusion equation is known to generate complex patterns (turing1990chemical) that can explain many natural processes. Previous work has successfully utilized this equation in the context of GNNs (haber2019imexnet; choi2023gread; eliasof2024graph) for solving graph related problems.